Historical data show that customers who download music from a popular Web service spend approximately ​$23 per​ month, with a standard deviation of ​$4. Assume the spending follows the normal probability distribution. Find the probability that a customer will spend at least ​$20 per month. How much​ (or more) do the top 6​% of customers​ spend?

What is the probability that a customer will spend at least ​$20 per​ month? _______​(Round to four decimal places as​ needed.)

How much do the top 6​% of customers​ spend? ​__________(Round to the nearest cent as​ needed.)
Answer #1

solution Given thot 23 و ( 2 $ ( ) م - (20 ) م -= 20 - 23 2) ولله ) (5,25 - z) م - ا = 2266 . - ا د ة 6,723 7234,ہ probabilitUsing standard normal table plz>z) = 67. I-p (zzz) = 0.06 P (zcz) 1 -0.06 0.94 p (za 1.56)=0.94 1.56 Z using formula Z-score